# PPA-Game: Characterizing and Learning Competitive Dynamics Among Online Content Creators

### v2

@@ 1 Introduction
To bridge the gap between theoretical models and the complex dynamics observed among online content creators, we introduce the Proportional Payoff Allocation Game (PPA-Game) as a novel model to capture the competitive dynamics of online content creators. Adapting from  (Ben-Porat et al., 2019b, 2020; Prasad et al., 2023), we assume a game with $N$ players (content creators) and $K$ resources (content topics). In this context, players need to decide which resources to select. Each resource has an associated total payoff (indicative of the exposure for each content topic), and individual players have specific weights (or quality metrics) assigned to each resource. When several players choose the same resource, its payoff is distributed proportionally according to player weights—a mechanism resonating with other game-theoretical frameworks  (Caragiannis and Voudouris, 2016; Kelly, 1997; Hoefer and Jiamjitrak, 2017; Anbarci et al., 2023).
@@ 3.1 Background on Game Theory
The concepts of $\epsilon$-Nash equilibrium and pure Nash equilibrium (PNE) have been well-established in the game theory literature (Nisan et al., 2007). To set the context for our study, consider a game with $N$ players. Let $\mathcal{S}$ represent the set of potential strategies for these players. Each player $j$ adopts a strategy $\pi_{j}\in\mathcal{S}$. The collective strategy of all players is represented as $\bm{\pi}=\{\pi_{j}\}_{j=1}^{N}$, while $(\pi^{\prime},\bm{\pi}_{-j})$ denotes a scenario where player $j$ deviates from his original strategy $\pi_{j}$ to another strategy $\pi^{\prime}\in\mathcal{S}$. We denote the payoff of player $j$ for a given strategy profile $\bm{\pi}$ as $\mathcal{U}_{j}(\bm{\pi})$. With these notations, the $\epsilon$-Nash equilibrium is defined as:
@@
A strategy profile $\bm{\pi}\in\mathcal{S}^{N}$ for a game defined by $\mathcal{G}=\left(\mathcal{S},\{\mathcal{U}_{j}\}_{j=1}^{N}\right)$ is termed an $\epsilon$-Nash equilibrium if, for every strategy $\pi^{\prime}\in\mathcal{S}$ and for each player $j\in[N]$, $\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j})\leq\mathcal{U}_{j}(\bm{\pi})+\epsilon$. If $\bm{\pi}$ achieves a 00-Nash equilibrium, it’s termed a pure Nash equilibrium.
@@ 3.2 Game Formulation
Consider a scenario with $K$ distinct resources labeled by indices $[K]=\{1,2,\dots,K\}$ and $N$ players, where $N,K\geq 2$. Each resource, denoted by $k$, possesses a total payoff $\mu_{k}$ in the range $0<\mu_{k}\leq 1$. Simultaneously, each player, represented as $j$, has a weight ranging between $0\leq w_{j,k}\leq 1$ for resource $k$. Every player can choose from a strategy set $[K]$, resulting in an overall strategy profile $\bm{\pi}\in[K]^{N}$. Crucially, if multiple players opt for the same resource, the total payoff from that resource is distributed proportionally based on the individual player weights. This can be formalized as:
@@
Given $K$ resources with associated payoffs $\bm{\mu}=\{\mu_{k}\}_{k=1}^{K}$ ($0<\mu_{k}\leq 1$) and $N$ players characterized by weights $\bm{W}=\{w_{j,k}:j\in[N],k\in[K]\}$ ($0\leq w_{j,k}\leq 1$), the Proportional Payoff Allocation Game (PPA-Game) is formulated as $\mathcal{G}=\left(\mathcal{S},\left\{\mathcal{U}_{j}(\bm{\pi})\right\}_{j=1}^{N}\right)$, where $\mathcal{S}=[K]$ and
$\mathcal{U}_{j}(\bm{\pi})=\begin{cases}\frac{\mu_{k}w_{j,k}}{\sum_{j^{\prime}=1}^{N}\mathbb{I}[\pi_{j^{\prime}}=k]w_{j^{\prime},k}}&\text{if }\sum_{j^{\prime}=1}^{N}\mathbb{I}[\pi_{j^{\prime}}=k]w_{j^{\prime},k}>0,\\ 0&\text{otherwise}.\end{cases}\$ (1)
@@
It’s worth noting that any resource $k$ without players having positive weights will remain unchosen. Consequently, such a resource can be disregarded. Based on this observation, we can assert that for each $k\in[K]$, there exists a $j\in[N]$ such that $w_{j,k}>0$, and the maximum weight across players is $1$, i.e., $\max_{j\in[N]}w_{j,k}=1$ for every $k\in[K]$. We then demonstrate that how the PPA-Game can be used to model competitive behaviors among content creators within online recommender systems:
@@ 3.3.1 High Probability of PNE Existence
We randomly select $N$ and $K$ from the range $[10,100]$ and sample resource payoffs and player weights from four distributions supported on $[0,1]$. Among $20,000,000$ randomly generated game configurations, a PNE fails to exist in only one instance.
@@ 3.3.2 Existence of PNE in Specific Scenarios
$\forall k\neq k^{\prime},\quad\frac{\mu_{k}}{\mu_{k^{\prime}}}>\frac{N_{0}}{\epsilon_{0}}\quad\text{or}\quad\frac{\mu_{k}^{\prime}}{\mu_{k}}>\frac{N_{0}}{\epsilon_{0}}.$ (2)
@@ 3.3.4 Price of Anarchy
$\mathrm{PoA}=\frac{\max_{\bm{\pi}\in[K]^{N}}W_{\mathcal{G}}(\bm{\pi})}{\min_{\bm{\pi}\in[K]^{N}:\bm{\pi}\text{ is a PNE}}W_{\mathcal{G}}(\bm{\pi})}.$ (3)
Theorem 3.3.
@@ 4.1 Multi-Player Multi-Armed Bandits Framework
Suppose there are $T$ rounds. At each round $t\in[T]$, each arm (resource) $k$ offers a stochastic reward $X_{k}(t)$, independently and identically distributed according to a distribution $\xi_{k}\in\Xi$, with expected value $\mu_{k}$, supported on $(0,1]$. We denote the set of all possible such distributions as $\Xi$. The vector of rewards from all arms at round $t$ is represented as $\bm{X}(t)=\{X_{k}(t)\}_{k=1}^{K}$. During each round $t\in[T]$, $N$ players select arms concurrently, based solely on their past observations. We use $\pi_{j}(t)$ to indicate the arm chosen by player $j$ and $\bm{\pi}(t)=\{\pi_{j}(t)\}_{j=1}^{N}$ to denote the strategy profile of all players at that round.
@@ 4.2 Evaluation metrics
$\operatorname{NonPNE}(T)=\sum_{t=1}^{T}\mathbb{I}[\bm{\pi}(t)\text{ is not a most efficient PNE of }\mathcal{G}].$ (5)
@@ 4.3.2 Learning and Exploitation Phase
Here, $\gamma_{j,k}$ is randomly sampled from a uniform distribution over the interval $[0,\Gamma]$. As proven in  Theorem 4.1, under certain conditions, $\mathcal{G}^{\prime}$ has a unique most efficient PNE with probability 1, which also serves as a most efficient PNE in the original game $\mathcal{G}$. Now we introduce the details of Learning PNE and Exploitation Subphases.
@@
Let $10<\epsilon<$ 1 be a constant. Define the functions $F(u)$ and $G(u)$ as follows:
$F(u)=-\frac{u}{(4+3\Gamma)N}+\frac{1}{3N}\quad\text{and}\quad G(u)=-\frac{u}{4+3\Gamma}+\frac{1}{3}.$ (7)
@@ 4.4.1 Quantifying the Challenge of Identifying the Most Efficient PNE
$\delta^{\text{NE}}$ $:\displaystyle=\min\big{\{}\mathcal{U}_{j}(\bm{\pi})-\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j})$ : (8)
@@
[…]
@@
Set $N=3$ and $K=2$. Here we suppose that each player has the same weights on the resources (as per Condition (2) in Theorem 3.2). When arms’ rewards are $[1,0.7]$ and players’ weights are $[1,0.8,0.4]$, we have $0.0260.052=\delta^{\text{NE}}>\delta^{\text{NoNE}}=$ 0.026. By contrast, when arms’ rewards are $[1,0.6]$ and players’ weights are $[1,2/3,4/9]$, we have $0.0680.040=\delta^{\text{NE}}<\delta^{\text{NoNE}}=$ 0.068. Details of the two cases can be found in Tables 1 and 2 in Appendix B.
@@
Suppose $\Delta<\delta/(4K)$, $\Gamma\leq\delta/(4N)$, and the PNE exists for the game $\mathcal{G}$ in Definition 3.2. Then the PNE exists for the game $\mathcal{G}^{\prime}$ (Eq. 6). In addition, the most efficient PNE in the game $\mathcal{G}^{\prime}$ is unique with probability $1$ and the equilibrium is a most efficient PNE in the original game $\mathcal{G}$.
@@ 5.2 Baselines and Hyper-parameters
  • TotalReward: Proposed by Bande and Veeravalli (2019), this method emphasizes a shareable reward case, aiming to maximize the collective reward for all players. In scenarios where $N<K$, the optimal strategy for each round is to have all players select the top $N$ arms. When $N\geq K$, every arm should be selected in every round. This strategy operates in an explore-then-commit manner. Specifically, during the initial $\alpha\log T$ rounds, the algorithm explores arms randomly. Thereafter, it commits to the optimal strategy in each round. The hyper-parameter $\alpha$ is chosen from the set $\{100,200,500,1000,2000\}$.
@@ A.1 The Non-existence of PNE in Certain Scenarios
$\bm{W}=\begin{pmatrix}0.1657&0&0.5755&0\\ 0&0.0314&0&0.5623\\ 0&0&0.4217&0.9517\\ 0&1&0.4057&0\\ 1&0&0&0\\ 0&0&1&0\\ 0&0&0&1\\ \end{pmatrix}.$
We find an improvement cycle for players $1$, $2$, $3$, and $4$ in this example, as depicted in Fig. 3. In this figure, the large circles labeled A, B, C, and D represent resources. The blue numbers between these resource circles signify the players and the sequence of their actions, while the numbers inside the larger circles indicate the respective rounds. At round $1$, player $1$ perceives that resource $B$ offers a more favorable payoff and hence shifts to resource $B$. This transition decreases the payoff for player $2$, prompting him to switch to resource $C$ by round $2$. The cycle continues, and by round $4$, player $4$ transitions from resource $D$ to $B$, which subsequently reduces the payoff for player $1$. This leads player $1$ to revert to resource $A$ by round $5$, and in turn, prompts player $2$ to shift back from resource $C$ to $B$ by round $6$, thus perpetuating the cycle.
@@ A.2 High Probability of PNE Existence
We experiment with cases where the values of $N$ and $K$ are large, using Monte Carlo sampling. We randomly select values for $N$ and $K$, uniformly from $[10,100]$. Moreover, for each player-resource pair $(j,k)$, with $j\in[N]$ and $k\in[K]$, weights $w_{j,k}$ and payoffs $\mu_{k}$ are sampled from four different types of distributions.
@@
  • Beta Distribution: We opt for a beta distribution with parameters $\alpha=\beta=0.5$. This distribution is skewed towards the extremes of 00 and $1$, representing scenarios where values are highly polarized.
@@ Appendix C Omitted Proofs
$\frac{\mu_{k}w_{j,k}}{Q_{k}(\ell)+w_{j,k}}=\mathcal{U}_{j}(\bm{\pi}(s^{\prime}))<\mathcal{U}_{j}(\bm{\pi}(s^{\prime}+1))\leq\frac{\mu_{k^{\prime}}w_{j,{k^{\prime}}}}{Q_{k^{\prime}}(\ell)+w_{j,k^{\prime}}},$
@@
$\frac{\mu_{k_{1}}}{Q_{k_{1}}(\ell)+\epsilon_{0}}<\frac{\mu_{k_{2}}}{Q_{k_{2}}(\ell)+\epsilon_{0}}<\dots<\frac{\mu_{k_{t+1}}}{Q_{k_{t+1}}(\ell)+\epsilon_{0}}.$
Since $k_{i}$ ($11\leq i\leq t+$ 1) can only take value in $\{1,2,\dots,t\}$, there must exists two indices $i_{1}\neq i_{2}$ such that $k_{i_{1}}=k_{i_{2}}$. This leads to a contradiction since Section C.1 suggests that $\forall 1\leq i_{1},i_{2}\leq t+1$, $k_{i_{1}}\neq k_{i_{2}}$. Now we can conclude that all improvement paths end in finite steps and the PNE exists. ∎
@@
Here $\Phi(\bm{\pi}(s))$ is a sorted vector and each element in the vector corresponds to a unit reward in a resource. Note that when $W_{k}(\bm{\pi}(s))=\sum_{j=1}^{N}\mathbb{I}[\pi_{j}(s)=k]w_{j}=0$ for some $s$ and $k$, we set the value $\mu_{k}/W_{k}(\bm{\pi}(s))$ be infinity ($\infty$). We suppose that $a<\infty$ for any real number $a$ and $\infty$=$\infty$ = ∞.
@@
$\frac{\mu_{\pi_{j}(s)}}{W_{\pi_{j}(s)}(\bm{\pi}(s+1))}\geq\frac{\mu_{\pi_{j}(s)}}{W_{\pi_{j}(s)}(\bm{\pi}(s))}\quad\text{and}\quad\frac{\mu_{\pi_{j}(s+1)}}{W_{\pi_{j}(s+1)}(\bm{\pi}(s))}>\frac{\mu_{\pi_{j}(s+1)}}{W_{\pi_{j}(s+1)}(\bm{\pi}(s+1))}.$
@@
$-\frac{W_{\pi_{j}(s+1)}\left(\bm{\pi}(s)\right)}{W_{\pi_{j}(s+1)}(\bm{\pi}(s+1))}=-\frac{W_{k(s)}\left(\bm{\pi}(s)\right)}{w_{j,k(s)}+W_{k(s)}\left(\bm{\pi}(s)\right)}>-\frac{W_{\pi_{j}(s)}(\bm{\pi}(s))-w_{j,\pi_{j}(s)}}{W_{\pi_{j}(s)}(\bm{\pi}(s))}=-\frac{W_{\pi_{j}(s)}(\bm{\pi}(s+1))}{W_{\pi_{j}(s)}(\bm{\pi}(s))}.$
@@
In the PPA-Game, the strategy profile that maximizes social welfare assigns all players to the top-$N$ resources if $N\leq K$, and if $N>K$, the top $K$ players select all $K$ resources. Consequently, the maximal social welfare is $\sum_{i=1}^{\min\{N,K\}}\mu_{i}$. Denote $W^{\max}=\sum_{i=1}^{\min\{N,K\}}\mu_{i}$ as the numerator of PoA given in Eq. 3.
Suppose $\bm{\pi}$ is a PNE. Let =$\mathcal{M}=\{1,2,\dots,\min\{N,K\}\}$ and $\mathcal{C}=\{j\in[K]:\exists i\in[N],\pi_{i}=j\}$. As a result, by the definition of social welfare and the utility function given in Eq. 1, we have that
@@
Note that since $\sum_{i=1}^{N}\mathcal{U}_{i}(\bm{\pi})=W_{\mathcal{G}}(\bm{\pi})$. There must exist one player $i^{*}$ such that $\mathcal{U}_{i}(\bm{\pi})\leq W_{\mathcal{G}}(\bm{\pi})/N$. Since $\bm{\pi}$ is a PNE, player $i^{*}$ will not benefit by deviation, which means that $\𝒞\forall j\in\mathcal{M}\backslash\mathcal{C}$, $\mu_{j}\leq W_{\mathcal{G}}(\bm{\pi})/N$. As a result,
@@
Note that we must have $1\in\mathcal{M}$ and $1\in\mathcal{C}$. Therefore, |\$|\mathcal{M}\backslash\mathcal{C}|\leq|\mathcal{M}|-1=\min\{N,K\}-1$. This gives us
@@
Because $\bm{\pi}$ is not a most efficient equilibrium of $\mathcal{G}$ by assumption, there must exist another equilibrium $\tilde{\bm{\pi}}$ of $\mathcal{G}$ such that $W_{\mathcal{G}}(\tilde{\bm{\pi}})>W_{\mathcal{G}}(\bm{\pi})$. By Lemma C.4, there must exist $\tilde{k}>k$ such that in the equilibrium $\tilde{\bm{\pi}}$, all the arms in $[\tilde{k}]$ is chosen by at least one player and all other arms are not chosen by any player. By part (a), we have that $\tilde{\bm{\pi}}$ is also an equilibrium of $\mathcal{G}^{\prime}$, and we have
@@
(c) We finally demonstrate that the most efficient Nash equilibrium of $\mathcal{G}^{\prime}$ is unique with probability $1$.
@@
$\,\mathbb{P}\left[\exists\bm{\pi}^{\prime},\bm{\pi}^{\prime\prime}\in[K]^{N},\bm{\pi}^{\prime}\text{ and }\bm{\pi}^{\prime\prime}\text{ are themost efficient Nash equilibria of }\mathcal{G}^{\prime},W_{\mathcal{G}^{\prime}}(\bm{\pi}^{\prime})=W_{\mathcal{G}^{\prime}}(\bm{\pi}^{\prime\prime})\right]$
@@
Note that this occurs with probability 00 because $\gamma_{j,k}$ are sampled uniformly from $[0,\Gamma]$ and the right-hand side of the above equation is a constant when $\bm{\pi}^{\prime}$ and $\bm{\pi}^{\prime\prime}$ are fixed. As a result, we have $\forall\bm{\pi}^{\prime},\bm{\pi}^{\prime\prime}\in[K]^{N}$, $\mathbb{P}[W_{\mathcal{G}^{\prime}}(\bm{\pi}^{\prime})=W_{\mathcal{G}^{\prime}}(\bm{\pi}^{\prime\prime})]=0$ and the left-hand side of Section C.4 is 00. Now the claim follows. ∎
@@
Suppose the conditions in Theorem 4.2 hold. Let $\bm{\pi}^{*}$ be the unique most efficient Nash equilibrium w.r.t. the game $\mathcal{G}^{\prime}$ according to Theorem 4.1. Then with probability $1$ the following holds. The state
$Z^{*}=\{(m_{j}^{*},a_{j}^{*},u_{j}^{*})\}_{j\in[N]}\text{ with }\forall j\in[N],m_{j}^{*}=C,a_{j}^{*}=\pi_{j}^{*},u_{j}^{*}=\mathcal{U}^{\prime}_{j}(\bm{\pi}^{*})$
is the unique stochastically stable state of the regular perturbed Markov process $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$. In addition, for any $20<\alpha<1/$ 2, there exists a small enough $\epsilon$ such that the probability of $Z^{*}$ in the stationary distribution $p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})$ is larger than $1/(2(1-\alpha))$.

### v3

@@ 1 Introduction
To bridge the gap between theoretical models and the complex dynamics observed among online content creators, we introduce the Proportional Payoff Allocation Game (PPA-Game) as a novel model to capture the competitive dynamics of online content creators. Adapting from  (Ben-Porat et al., 2019b, 2020; Prasad et al., 2023), we assume a game with NN players (content creators) and $K$ resources (content topics). In this context, players need to decide which resources to select. Each resource has an associated total payoff (indicative of the exposure for each content topic), and individual players have specific weights (or quality metrics) assigned to each resource. When several players choose the same resource, its payoff is distributed proportionally according to player weights—a mechanism resonating with other game-theoretical frameworks  (Caragiannis and Voudouris, 2016; Kelly, 1997; Hoefer and Jiamjitrak, 2017; Anbarci et al., 2023).
@@ 3.1 Background on Game Theory
The concepts of $\epsilon$-Nash equilibrium and pure Nash equilibrium (PNE) have been well-established in the game theory literature (Nisan et al., 2007). To set the context for our study, consider a game with NN players. Let $\mathcal{S}$ represent the set of potential strategies for these players. Each player $j$ adopts a strategy $\pi_{j}\in\mathcal{S}$. The collective strategy of all players is represented as $\bm{\pi}=\{\pi_{j}\}_{j=1}^{N}$, while $(\pi^{\prime},\bm{\pi}_{-j})$ denotes a scenario where player $j$ deviates from his original strategy $\pi_{j}$ to another strategy $\pi^{\prime}\in\mathcal{S}$. We denote the payoff of player $j$ for a given strategy profile $\bm{\pi}$ as $\mathcal{U}_{j}(\bm{\pi})$. With these notations, the $\epsilon$-Nash equilibrium is defined as:
@@
A strategy profile $\bm{\pi}\in\mathcal{S}^{N}$ for a game defined by $\mathcal{G}=\left(\mathcal{S},\{\mathcal{U}_{j}\}_{j=1}^{N}\right)$ is termed an $\epsilon$-Nash equilibrium if, for every strategy $\pi^{\prime}\in\mathcal{S}$ and for each player $j\in[N]$, $\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j})\leq\mathcal{U}_{j}(\bm{\pi})+\epsilon$. If $\bm{\pi}$ achieves a 0-Nash equilibrium, it’s termed a pure Nash equilibrium.
@@ 3.2 Game Formulation
Consider a scenario with $K$ distinct resources labeled by indices $K]=\{1,2,\dots,K\}$ and NN players, where $N,K\geq 2$. Each resource, denoted by $k$, possesses a total payoff $\mu_{k}$ in the range $0<\mu_{k}\leq 1$. Simultaneously, each player, represented as $j$, has a weight ranging between $0\leq w_{j,k}\leq 1$ for resource $k$. Every player can choose from a strategy set $[K]$, resulting in an overall strategy profile $\bm{\pi}\in[K]^{N}$. Crucially, if multiple players opt for the same resource, the total payoff from that resource is distributed proportionally based on the individual player weights. This can be formalized as:
@@
Given $K$ resources with associated payoffs $\bm{\mu}=\{\mu_{k}\}_{k=1}^{K}$ ($0<\mu_{k}\leq 1$) and NN players characterized by weights $\bm{W}=\{w_{j,k}:j\in[N],k\in[K]\}$ ($0\leq w_{j,k}\leq 1$), the Proportional Payoff Allocation Game (PPA-Game) is formulated as $\mathcal{G}=\left(\mathcal{S},\left\{\mathcal{U}_{j}(\bm{\pi})\right\}_{j=1}^{N}\right)$, where $\mathcal{S}=[K]$ and
[…]
@@
It’s worth noting that any resource $k$ without players having positive weights will remain unchosen. Consequently, such a resource can be disregarded. Based on this observation, we can assert that for each $k\in[K]$, there exists a $j\in[N]$ such that $w_{j,k}>0$, and the maximum weight across players is 11, i.e., $\max_{j\in[N]}w_{j,k}=1$ for every $k\in[K]$. We then demonstrate that how the PPA-Game can be used to model competitive behaviors among content creators within online recommender systems:
@@ 3.3.1 High Probability of PNE Existence
We randomly select NN and $K$ from the range $[10,100]$ and sample resource payoffs and player weights from four distributions supported on $[0,1]$. Among $20,000,000$ randomly generated game configurations, a PNE fails to exist in only one instance.
@@ 3.3.2 Existence of PNE in Specific Scenarios
[…]
@@ 3.3.4 Price of Anarchy
[…]
Theorem $3.$
@@ 4.1 Multi-Player Multi-Armed Bandits Framework
Suppose there are $T$ rounds. At each round $t\in[T]$, each arm (resource) $k$ offers a stochastic reward $X_{k}(t)$, independently and identically distributed according to a distribution $\xi_{k}\in\Xi$, with expected value $\mu_{k}$, supported on $(0,1]$. We denote the set of all possible such distributions as $\Xi$. The vector of rewards from all arms at round $t$ is represented as $\bm{X}(t)=\{X_{k}(t)\}_{k=1}^{K}$. During each round $t\in[T]$, NN players select arms concurrently, based solely on their past observations. We use $\pi_{j}(t)$ to indicate the arm chosen by player $j$ and $\bm{\pi}(t)=\{\pi_{j}(t)\}_{j=1}^{N}$ to denote the strategy profile of all players at that round.
@@ 4.2 Evaluation metrics
[…]
@@ 4.3.2 Learning and Exploitation Phase
Here, $\gamma_{j,k}$ is randomly sampled from a uniform distribution over the interval $0,\Gamma]$. As proven in  Theorem 4.1, under certain conditions, $\mathcal{G}^{\prime}$ has a unique most efficient PNE with probability 1, which also serves as a most efficient PNE in the original game $\mathcal{G}$. Now we introduce the details of Learning PNE and Exploitation Subphases.
@@
Let $0<\epsilon<1$ be a constant. Define the functions $F(u)$ and $G(u)$ as follows:
[…]
@@ 4.4.1 Quantifying the Challenge of Identifying the Most Efficient PNE
$\delta^{\text{NE}}$ $=\min\big\{\mathcal{U}_{j}(\bm{\pi})-\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j}):$ (8)
@@
$\delta^{\text{NoNE}}$ $=\sup\left\{\delta^{\prime}:\forall\bm{\pi}\in A^{\text{NoNE}},\bm{\pi}\text{ is not a }\delta^{\prime}\text{-Nash equilibrium}\right\}.$
@@
Set $N=3$ and $K=2$. Here we suppose that each player has the same weights on the resources (as per Condition (2) in Theorem 3.2). When arms’ rewards are $[1,0.7]$ and players’ weights are $[1,0.8,0.4]$, we have $0.052=\delta^{\text{NE}}>\delta^{\text{NoNE}}=0.026$. By contrast, when arms’ rewards are $[1,0.6]$ and players’ weights are $[1,2/3,4/9]$, we have $0.040=\delta^{\text{NE}}<\delta^{\text{NoNE}}=0.068$. Details of the two cases can be found in Tables 1 and 2 in Appendix B.
@@
Suppose $\Delta<\delta/(4K)$, $\Gamma\leq\delta/(4N)$, and the PNE exists for the game $\mathcal{G}$ in Definition 3.2. Then the PNE exists for the game $\mathcal{G}^{\prime}$ (Eq. 6). In addition, the most efficient PNE in the game $\mathcal{G}^{\prime}$ is unique with probability 11 and the equilibrium is a most efficient PNE in the original game $\mathcal{G}$.
@@ 5.2 Baselines and Hyper-parameters
  • TotalReward: Proposed by Bande and Veeravalli (2019), this method emphasizes a shareable reward case, aiming to maximize the collective reward for all players. In scenarios where $N<K$, the optimal strategy for each round is to have all players select the top NN arms. When $N\geq K$, every arm should be selected in every round. This strategy operates in an explore-then-commit manner. Specifically, during the initial $\alpha\log T$ rounds, the algorithm explores arms randomly. Thereafter, it commits to the optimal strategy in each round. The hyper-parameter $\alpha$ is chosen from the set $\{100,200,500,1000,2000\}$.
@@ A.1 The Non-existence of PNE in Certain Scenarios
[…]
We find an improvement cycle for players 11, 22, 33, and 44 in this example, as depicted in Fig. 3. In this figure, the large circles labeled A, B, C, and D represent resources. The blue numbers between these resource circles signify the players and the sequence of their actions, while the numbers inside the larger circles indicate the respective rounds. At round 11, player 11 perceives that resource $B$ offers a more favorable payoff and hence shifts to resource $B$. This transition decreases the payoff for player 22, prompting him to switch to resource $C$ by round 22. The cycle continues, and by round 44, player 44 transitions from resource $D$ to $B$, which subsequently reduces the payoff for player 11. This leads player 11 to revert to resource AA by round 55, and in turn, prompts player 22 to shift back from resource $C$ to $B$ by round 66, thus perpetuating the cycle.
@@ A.2 High Probability of PNE Existence
We experiment with cases where the values of NN and $K$ are large, using Monte Carlo sampling. We randomly select values for NN and $K$, uniformly from $[10,100]$. Moreover, for each player-resource pair $(j,k)$, with $j\in[N]$ and $k\in[K]$, weights $w_{j,k}$ and payoffs $\mu_{k}$ are sampled from four different types of distributions.
@@
  • Beta Distribution: We opt for a beta distribution with parameters $\alpha=\beta=0.5$. This distribution is skewed towards the extremes of 0 and 11, representing scenarios where values are highly polarized.
@@ Appendix C Omitted Proofs
[…]
@@
[…]
Since $k_{i}$ ($1\leq i\leq t+1$) can only take value in $\{1,2,\dots,t\}$, there must exists two indices $i_{1}\neq i_{2}$ such that $k_{i_{1}}=k_{i_{2}}$. This leads to a contradiction since Section C.1 suggests that $\forall 1\leq i_{1},i_{2}\leq t+1$, $k_{i_{1}}\neq k_{i_{2}}$. Now we can conclude that all improvement paths end in finite steps and the PNE exists. ∎
@@
Here $\Phi(\bm{\pi}(s))$ is a sorted vector and each element in the vector corresponds to a unit reward in a resource. Note that when $W_{k}(\bm{\pi}(s))=\sum_{j=1}^{N}\mathbb{I}[\pi_{j}(s)=k]w_{j}=0$ for some $s$ and $k$, we set the value $\mu_{k}/W_{k}(\bm{\pi}(s))$ be infinity ($\infty$). We suppose that $a<\infty$ for any real number $a$ and $\infty=\infty$.
@@
[…]
@@
$\frac{W_{\pi_{j}(s+1)}\left(\bm{\pi}(s)\right)}{W_{\pi_{j}(s+1)}(\bm{\pi}(s+1))}=-\frac{W_{k(s)}\left(\bm{\pi}(s)\right)}{w_{j,k(s)}+W_{k(s)}\left(\bm{\pi}(s)\right)}>-\frac{W_{\pi_{j}(s)}(\bm{\pi}(s))-w_{j,\pi_{j}(s)}}{W_{\pi_{j}(s)}(\bm{\pi}(s))}=-\frac{W_{\pi_{j}(s)}(\bm{\pi}(s+1))}{W_{\pi_{j}(s)}(\bm{\pi}(s))}.$
@@
In the PPA-Game, the strategy profile that maximizes social welfare assigns all players to the top-NN resources if $N\leq K$, and if $N>K$, the top $K$ players select all $K$ resources. Consequently, the maximal social welfare is $\sum_{i=1}^{\min\{N,K\}}\mu_{i}$. Denote $W^{\max}=\sum_{i=1}^{\min\{N,K\}}\mu_{i}$ as the numerator of PoA given in Eq. 3.
Suppose $\bm{\pi}$ is a PNE. Let $\mathcal{M}=\{1,2,\dots,\min\{N,K\}\}$ and $\mathcal{C}=\{j\in[K]:\exists i\in[N],\pi_{i}=j\}$. As a result, by the definition of social welfare and the utility function given in Eq. 1, we have that
@@
Note that since $\sum_{i=1}^{N}\mathcal{U}_{i}(\bm{\pi})=W_{\mathcal{G}}(\bm{\pi})$. There must exist one player $i^{*}$ such that $\mathcal{U}_{i}(\bm{\pi})\leq W_{\mathcal{G}}(\bm{\pi})/N$. Since $\bm{\pi}$ is a PNE, player $i^{*}$ will not benefit by deviation, which means that $\forall j\in\mathcal{M}\backslash\mathcal{C}$, $\mu_{j}\leq W_{\mathcal{G}}(\bm{\pi})/N$. As a result,
@@
Note that we must have $1\in\mathcal{M}$ and $1\in\mathcal{C}$. Therefore, $|\mathcal{M}\backslash\mathcal{C}|\leq|\mathcal{M}|-1=\min\{N,K\}-1$. This gives us
@@
Because $\bm{\pi}$ is not a most efficient equilibrium of $\mathcal{G}$ by assumption, there must exist another equilibrium $\tilde{\bm{\pi}}$ of $\mathcal{G}$ such that $W_{\mathcal{G}}(\tilde{\bm{\pi}})>W_{\mathcal{G}}(\bm{\pi})$. By Lemma C.4, there must exist $\tilde{k}>k$ such that in the equilibrium $\tilde{\bm{\pi}}$, all the arms in $\tilde{k}$] is chosen by at least one player and all other arms are not chosen by any player. By part (a), we have that $\tilde{\bm{\pi}}$ is also an equilibrium of $\mathcal{G}^{\prime}$, and we have
@@
(c) We finally demonstrate that the most efficient Nash equilibrium of $\mathcal{G}^{\prime}$ is unique with probability 11.
@@
[…]
@@
Note that this occurs with probability 0 because $\gamma_{j,k}$ are sampled uniformly from $0,\Gamma]$ and the right-hand side of the above equation is a constant when $\bm{\pi}^{\prime}$ and $\bm{\pi}^{\prime\prime}$ are fixed. As a result, we have $\forall\bm{\pi}^{\prime},\bm{\pi}^{\prime\prime}\in[K]^{N}$, $\mathbb{P}[W_{\mathcal{G}^{\prime}}(\bm{\pi}^{\prime})=W_{\mathcal{G}^{\prime}}(\bm{\pi}^{\prime\prime})]=0$ and the left-hand side of Section C.4 is 0. Now the claim follows. ∎
@@
Suppose the conditions in Theorem 4.2 hold. Let $\bm{\pi}^{*}$ be the unique most efficient Nash equilibrium w.r.t. the game $\mathcal{G}^{\prime}$ according to Theorem 4.1. Then with probability 11 the following holds. The state
[…]
is the unique stochastically stable state of the regular perturbed Markov process $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$. In addition, for any $0<\alpha<1/2$, there exists a small enough $\epsilon$ such that the probability of $Z^{*}$ in the stationary distribution $p_{Z^{*}}(\epsilon;\mathcal{G}^{\prime})$ is larger than $1/(2(1-\alpha))$.